If y varies directly as x^3 and y =400 when x=10 ,
a. Find y in terms of x.
b. Find y when x=5
c.Find x when y=25 3/5

Answer by ewatrrr(24785)


Hi
If y varies directly as x^3 and y =400 when x=10 400 = k*1000 k = 4/10 = 2/5
a. Find y in terms of x: y = (2/5)x^3
b. Find y when x=5 y = (2/5)125 = 50
c.Find x when y=25 3/5 128/5 = 2/5x^3 x^3 = 64 x = 4